rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_blue
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_green
icon_orange
 

FS#2942 — FS#6902 — upgrade sup2T Londres - 21/06

Attached to Project— Network
Maintenance
the whole network
CLOSED
100%
We intervene this night (from 20 to 21/06) on London to migrate our two routers ldn-1 et ldn-5 to sup2T. The routers will be isolated (successively!) from the network the time of rhe maintenance. The raffic to UK and US/CA may be impacted during work. We start at 2:00 CEST.
Date:  Tuesday, 26 June 2012, 21:37PM
Reason for closing:  Done
Comment by OVH - Thursday, 21 June 2012, 03:09AM

We are starting the work tasks on ldn-5-6k. We are isolating the router of the backbone.


Comment by OVH - Thursday, 21 June 2012, 03:12AM

We have switched off the router and we are replacing the cards of the chassis.


Comment by OVH - Thursday, 21 June 2012, 03:16AM

The router is up on sup2T. All links are reconnected. ldn-5-6k is performing properly. We are going to isolate ldn-1-6k.


Comment by OVH - Thursday, 21 June 2012, 04:08AM

ldn-1-6k is cut. We are upgrading sup2T.


Comment by OVH - Thursday, 21 June 2012, 08:53AM

Thereafter the wiring error on the site, the fiber dwdm to Roubaix has been cut and routing to ldn-5 (UK traffic) the US part of the interrupted backbone.


Comment by OVH - Thursday, 21 June 2012, 08:59AM

We had some routing problems which has impacted the traffic between 04:00 and 06:00. BGP sessions to the route-reflectors are not mounted correctly on the new routers in sup2T, probably due to the incompatibility. We have updated one of the RF which has fixed the problem (https://status.ovh.net/?do=details&id=2946). New routers are in production.